OpenStudy (anonymous):

\[(x^{2}-5x+7)^{x^2-1}=1\]

OpenStudy (anonymous):

I found x=1 and x=-1, but there are supposed to be two more answers, x=2 and x=3. How do you find them?

OpenStudy (anonymous):

You got two cases either : x^2-5x+7=1 or x^2-1=0

OpenStudy (anonymous):

x1=2; x2=3; x3=-1; x4=1;

OpenStudy (anonymous):

You're right, the "base=1" case completely slipped my mind.

OpenStudy (cwrw238):

take logs of both sides (x^1-1) ln(x^2- 5x + 7) = 0 x^2 -1 = 0 gives you the +1 and -1 roots ln(x^2 - 5x + 7) = 0 so x^2 - 5x + 7 = 1 will give you the other 2 roots

OpenStudy (anonymous):

good one @cwrw238

OpenStudy (anonymous):

Thank you both.